find an equation, in slope intercept form, that passes through point (-4, -1) with slope -2.
Answer by mananth(16946)

m= -2

Plug value of the slope and point ( -4 , -1 ) in
Y = m x + b
-1.00 = 8 + b
b= -1 - 8
b= -9
So the equation will be
Y = -2 x -9
m.ananth@hotmail.ca